OpenStudy (anonymous):

PLEASE HELP!!! Joselyn is a manager at a sign-painting company. She has two painters, Allen and Brianne. Allen can complete a large project in 16 hours. Brianne can complete the project in 18 hours. Joselyn wants to know how long it will take them to complete the project together. Write an equation and solve for the time it takes Allen and Brianne to complete the project together.

OpenStudy (anonymous):

@nincompoop @hartnn

hartnn (hartnn):

Let them take 'x' hours to complete the work by working together.

hartnn (hartnn):

` Allen can complete a large project in 16 hours` if Allen works alone, how much of the work she can complete in 1 hour?

OpenStudy (anonymous):

1/16 of the project?

hartnn (hartnn):

yes! what about Brianne?

OpenStudy (anonymous):

1/18

hartnn (hartnn):

yes, what if they work together?

OpenStudy (anonymous):

um, 1/16 + 1/18

OpenStudy (anonymous):

?

hartnn (hartnn):

good! now we know they both can do the entire work in 'x' hours by working together, we can say that they both will complete 1/x of the work in one hour. makes sense?

OpenStudy (anonymous):

yes

hartnn (hartnn):

and we already know they both can complete it in 1/16 +1/18 of the work in 1 hr. so we can equate: \(\Large \dfrac{1}{16} + \dfrac{1}{18} =\dfrac{1}{x}\) can you isolate x from here?

OpenStudy (anonymous):

um 1/x = 17/144 ?

hartnn (hartnn):

bdw, this is the equation they want : \(\Large \dfrac{1}{16} + \dfrac{1}{18} =\dfrac{1}{x}\) now you're solving for the time, yes, 1/x= 17/144 so, x = 144/7

hartnn (hartnn):

x= 144/17**

OpenStudy (anonymous):

ok. what's next?

hartnn (hartnn):

it will take them 144/17 or 8.47 hrs if they work together. and thats it!

OpenStudy (anonymous):

ok thank you :)

hartnn (hartnn):

welcome ^_^